OpenStudy (anonymous):

simplify x^3 times x^-5

OpenStudy (anonymous):

Add the exponents.

OpenStudy (anonymous):

= x^3 + (-5) = x^-2 = 1/x^2

OpenStudy (saifoo.khan):

\[\Huge a^ b \times a^c = a^{b+c}\]

OpenStudy (anonymous):

3x^2 * -5x^-4

OpenStudy (saifoo.khan):

@maryamali , try posting the question on left pane! :)

OpenStudy (saifoo.khan):

in*

OpenStudy (anonymous):

3*-5*x^2 + (-4) = -15x^-2 = -15/x^2

OpenStudy (anonymous):

hey harikat, i don't know if i'm wrong or you are, is it meant to be differentiated?

OpenStudy (saifoo.khan):

No, we have to multiply. \[\large (3 * -5) (x^2 * -x^4)\to (-15)(x^{-2})\to -\frac{15}{x^2}\]

OpenStudy (anonymous):

why differentiated????

OpenStudy (anonymous):

don't you have to add the powers? so it would be -2?

OpenStudy (anonymous):

the solution is as I and saifoo have shown

OpenStudy (anonymous):

ok thanku